OpenStudy (anonymous):

simplify: sqrt(200) - sqrt(32)

OpenStudy (anonymous):

200 = 100*2 32 = 16*2 can you simplify now?

OpenStudy (anonymous):

200 and 32 ?

OpenStudy (anonymous):

i factored the radicands so you can simplify easier....

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow \sqrt{200} = \sqrt{2 * 2 * 2 * 5 * 2 * 2 * 5} }\) \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{200} = 20\sqrt{2} }\) AND \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{32} = \sqrt{2 * 2 * 2 * 2 * 2} }\) \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{32} = 2\sqrt{2} }\) Imagine sqrt2 as a. What will be 20a - 2a?

OpenStudy (anonymous):